# First Demonstration of Underground Muon Imaging at an Archaeological Site in Ancient Jerusalem

### v1

@@ III.3 Measured Angular Ground Depth
By applying the measured differential flux to lookup tables from muon transport simulations, we derive the angular ground depth $D(\tau_{x},\tau_{y})$, which acts as a proxy for the integrated density of the overburden. Figure 7 shows the resulting density map. This map, the primary result of our study, reveals variations in the overburden that may indicate subsurface anomalies such as voids or high-density inclusions.
Figure 7: The angular ground depth (meters) $D(\tau_{x},\tau_{y})$ derived from the transformation of the observed muon flux.
@@ IV Discussion
The measurement demonstrates that muons can be used to map underground features. The discrepancies between the measured and expected muon flux and overburden can be attributed to several factors. First, the detector efficiency is low (less than 20%), primarily due to a suboptimal triggering scheme and excessively high thresholds for muon hit identification. This inefficiency results in significant undercounting of muon events and a reduced flux relative to simulation predictions. Optimization of the trigger logic and recalibration of the thresholds are necessary steps for improving the measurement’s sensitivity and accuracy.
Second, the simulation of the atmospheric muon flux assumed a sea-level model. Since Jerusalem is located approximately 650 m above sea level, the actual muon flux is higher, with a slightly altered angular distribution. This altitude effect introduces a systematic discrepancy in both the overall muon rate and the angular profile, affecting the accuracy of the derived overburden.
Third, while the interior part of the cistern was precisely scanned with LIDAR, the ground level was assumed to be flat and without substantial infrastructure. However, this approximation is inaccurate; the ground above Jeremiah’s cistern is not flat and has several $\mathcal{O}(meter)$ features, namely massive stone walls, large olive trees, and buildings surrounding the premise. These were not incorporated into the simulations.
Additional uncertainties stem from limitations in the muon transport simulations, environmental variations, and potential systematic uncertainties in detector calibration. The current effort focuses on incorporating altitude corrections into the simulation models, improving calibration procedures, and deploying multiple detectors to achieve three-dimensional subsurface imaging.
@@ V Conclusion
This study has successfully demonstrated the utility of muon imaging as an effective non-invasive imaging technique for mapping underground features, specifically within the archaeological context of Jeremiah’s cistern in the City of David, Jerusalem. We could derive key insights into the subsurface density variation and overburden characteristics by deploying a custom muon detector and integrating advanced LIDAR mapping techniques. While discrepancies between the measured and expected muon flux were identified, primarily attributed to detector efficiency, altitude effects, and the inherent complexities of the ground structure, the findings pave the way for further refinement and optimization of the muon detection process. Follow-up work will address these uncertainties by enhancing calibration methods, incorporating altitude adjustments in simulations, and expanding the network of detectors for comprehensive three-dimensional imaging. Overall, this research highlights the potential of muon tomography in archaeological investigations and encourages its continued application in exploring buried structures and understanding ancient civilizations.
@@ VI Data Availability
The data supporting this study’s findings are available from the corresponding author upon reasonable request.
@@ Acknowledgements.
This work was supported in part by the Pazy and Cogito foundations. The site is located in the City of David in the Jerusalem Walls National Park. Y. Shalev and F. Vukosavović excavated on behalf of the IAA, Permit No. 8928.

### v2

@@ III.3 Measured Angular Ground Depth
By applying the lookup tables from muon transport simulations to the measured data, we derive the angular ground depth $D(\tau_{x},\tau_{y})$, which acts as a proxy for the integrated density of the overburden. Fig. 9 shows the resulting density map. This map, the primary result of our study, reveals variations in the overburden that may indicate subsurface anomalies, such as voids or high-density inclusions.
Figure 9: The angular ground depth $D(\tau_{x},\tau_{y})$ derived by applying the lookup tables from muon transport simulations to the measured data.
@@ IV Discussion
The measurement shows that muons can map underground features and resolve structural variations in the overburden. The angular muon flux distribution (Fig. 8) highlights anisotropies consistent with the ventilation shafts seen in the LiDAR scan (Fig. 7), which was integrated into our simulation framework. The match between simulated and measured angular patterns confirms that the features correspond to actual geometric structures above the detector. However, while the relative angular trends are meaningful, the overall flux normalization is affected by several factors: detector efficiency, the assumption of uniform overburden density, and the flat surface profile approximation above the cistern, all of which differ from actual site conditions. These discrepancies are discussed below.
First, the detector efficiency is low (less than 20%), mainly because of a suboptimal trigger scheme and excessively high detection thresholds. This inefficiency leads to a significant under-counting of muon events and a reduced flux compared to simulation predictions. Improving the trigger logic and recalibrating the thresholds are essential steps to enhance the measurement’s sensitivity and accuracy.
Second, the simulation of the atmospheric muon flux assumed a sea-level model. Since Jerusalem is approximately 650 m above sea level, the actual muon flux is higher, with a slightly different angular distribution. Empirical models of cosmic-ray muon production Reyna (2006); Bugaev et al. (1998) indicate that at this elevation, the vertical muon flux increases by approximately 15–20% compared to sea level. This increase results from the thinner atmosphere, which allows more muons to reach the surface. Not applying altitude correction causes a systematic difference in both the absolute muon rate and its angular distribution, thus impacting the inferred overburden.
Third, although the cistern’s interior was accurately scanned with LiDAR, the simulations assumed a flat ground surface above it. In reality, the terrain has significant features on the meter scale, such as large stone walls, mature olive trees, and nearby buildings. These were not included in the simulation geometry, and their omission adds extra error in modeling muon attenuation.
Finally, limitations in muon transport simulations also contribute to the observed discrepancies. These simulations, based on the energy loss equation (Eq. 8), incorporate ionization and radiative losses via energy-dependent coefficients $a(E)$ and $b(E)$. The lookup tables derived from these models assume a homogeneous material and simplified geometry, neglecting local inhomogeneities or complex surface structures. This simplification can cause local errors in the reconstructed density map. Additional uncertainties come from environmental variations and potential systematic biases in detector calibration. Ongoing efforts include adding altitude corrections, improving calibration procedures, and expanding the detector array to enable full three-dimensional subsurface imaging.
Furthermore, as part of our forward-looking strategy, we plan to incorporate AI-based image processing techniques to reduce noise, improve flux map reconstruction, and detect subtle anomalies in challenging data conditions. These efforts will build on advances in other fields such as medical tomography and geophysical imaging, and we believe they have strong potential for archaeological muon imaging as well.
@@ V Conclusion
This study demonstrates the utility of muon imaging as an effective, non-invasive technique for mapping underground features, specifically in the archaeological context of the cistern in the City of David, Jerusalem. We could derive key insights into the subsurface density variation and overburden characteristics by deploying a custom muon detector and integrating high-resolution LiDAR mapping. While discrepancies between the measured and expected muon flux were identified, primarily attributed to detector efficiency, altitude effects, and the inherent complexity of the overlying structures, these findings pave the way for further refinement and optimization of the muon detection process. Future work will address these uncertainties by improving calibration methods, incorporating altitude adjustments in simulations, and expanding the detector network to enable comprehensive three-dimensional imaging. Overall, this study highlights the potential of muon tomography for archaeological investigations and supports its continued application in exploring buried structures and understanding ancient civilizations.
@@ VI Data Availability
The data supporting the findings of this study are available from the corresponding author upon reasonable request. Due to the technical nature of the data and its dependence on specific analysis procedures, effective use requires further explanation from the authors.
@@ Acknowledgements.
This work was supported in part by the Pazy and Cogito foundations. The site is located in the City of David within the Jerusalem Walls National Park. Y. Shalev and F. Vukosavović excavated on behalf of the IAA, Permit No. 8928.